Find the general solution of y'' + y'-6y=(9x-2)e^(2x). (Use the method of undetermined coefficients)

2. Find a general solution of y + y - 6y = (9.C -- 2)e2.. (22 pts, use the method of undeter- mined coefficients.)

Please show all work and steps!
Answer #1

Q:. Given Diff Equ is .. +y! Loy - (94-2) ed? --O Auxiliary Equation is! 02 +D- D ID+31 ) D= 2,-3 c.f. = cient cge 3 - we onom from o 4 ③ [lonix +(2 A1 +502)) en 194-2 je24 - 1o A1 = 9 & 2 A1+ 5 Ag ang A1 = 17 + 27 +54233 SA2: -2-3 SA2 -- I. Yp - E